# Deep neural networks for choice analysis: Enhancing behavioral regularity with gradient regularization

### v1

@@ Appendix C Training and validation performance
mean (std.) No GR PGR UGR LGR No GR PGR UGR LGR MNL
Log-likelihood $-633.6$ $\mathbf{-633.0}$ $-634.1$ $-634.2$ $\underline{-683.0}$ $-683.1$ $-683.1$ $-683.1$693.8
(2.507) (2.474) (3.594) (2.794) (2.442) (2.470) (2.482) (2.467) (0)
Accuracy 0.739 0.739 0.739 0.739 0.726 0.727 0.727 0.727 0.737
(0.004) (0.004) (0.003) (0.003) (0.004) (0.003) (0.003) (0.003) (0)
$F_{1}$ score 0.702 0.710 0.705 0.708 0.668 0.669 0.670 0.669 0.690
(0.005) (0.004) (0.005) (0.003) (0.007) (0.006) (0.006) (0.006) (0)
Strong regularity 0.879 0.981 0.973 0.982 0.997 0.999 0.998 0.998 1.000
(0.058) (0.005) (0.014) (0.005) (0.005) (0.002) (0.003) (0.002) (0)
Weak regularity 0.928 0.999 0.997 0.999 0.998 0.999 0.999 0.999 1.000
(0.057) (0.001) (0.004) (0.001) (0.003) (0.002) (0.002) (0.002) (0)
Panel 2: Chicago dataset, norm-XGR
Log-likelihood $-633.6$ $\mathbf{-633.3}$ $-635.9$ $\mathbf{-633.3}$ $-683.0$ $-683.0$ $-687.4$ $\underline{-682.9}$ $-693.8$
(2.507) (2.414) (2.676) (2.989) (2.442) (2.421) (2.236) (2.366) (0)
Accuracy 0.739 0.741 0.741 0.741 0.726 0.725 0.720 0.725 0.737
(0.004) (0.004) (0.003) (0.003) (0.004) (0.004) (0.001) (0.004) (0)
$F_{1}$ score 0.702 0.702 0.696 0.699 0.668 0.666 0.655 0.665 0.690
(0.005) (0.005) (0.006) (0.005) (0.007) (0.007) (0.003) (0.006) (0)
Strong regularity 0.879 0.846 0.701 0.803 0.997 0.996 0.924 0.995 1.000
(0.058) (0.058) (0.044) (0.061) (0.005) (0.005) (0.03) (0.007) (0)
Weak regularity 0.928 0.900 0.766 0.858 0.998 0.998 0.935 0.997 1.000
(0.057) (0.057) (0.046) (0.062) (0.003) (0.003) (0.026) (0.006) (0)
Table C.2 (continued).
Panel 3: London dataset, sum-XGR
@@
mean (std.) No GR PGR UGR LGR No GR PGR UGR LGR MNL
Log-likelihood $-645.5$ $\mathbf{-642.8}$ $-644.2$ $-647.8$ $\underline{-638.8}$ $-639.9$ $-645.3$ $-639.9$ $-670.0$
(4.399) (5.709) (4.345) (3.915) (1.949) (2.182) (3.254) (1.985) (0)
Accuracy 0.721 0.722 0.725 0.723 0.729 0.726 0.726 0.726 0.735
(0.002) (0.003) (0.002) (0.005) (0.003) (0.002) (0.002) (0.002) (0)
$F_{1}$ score 0.719 0.721 0.722 0.718 0.723 0.720 0.719 0.720 0.730
(0.003) (0.003) (0.002) (0.004) (0.003) (0.002) (0.003) (0.002) (0)
Strong regularity 0.947 0.992 0.991 0.996 0.924 0.953 0.982 0.964 0.991
(0.032) (0.005) (0.012) (0.004) (0.026) (0.014) (0.011) (0.014) (0)
Weak regularity 0.971 0.999 0.997 1.000 0.957 1.000 1.000 1.000 1.000
(0.025) (0.001) (0.007) (0.001) (0.023) (0.001) (0.000) (0.001) (0)
Small sample scenario
Table C.3: Model performance in the training sets (1K-Random).
Panel 1: Chicago dataset, sum-XGR
@@
mean (std.) No GR PGR UGR LGR No GR PGR UGR LGR MNL
Log-likelihood $\mathbf{-494.8}$ $-511.0$ $-546.4$ $-526.2$ $-555.6$ $-580.2$ $-570.9$ $-583.1$ $-548.1$
(6.074) (8.266) (9.391) (8.871) (4.071) (6.789) (5.316) (9.196) (0)
Accuracy 0.740 0.732 0.691 0.718 0.711 0.691 0.696 0.690 0.715
(0.005) (0.011) (0.008) (0.01) (0.006) (0.007) (0.008) (0.007) (0)
$F_{1}$ score 0.690 0.682 0.578 0.648 0.634 0.575 0.588 0.573 0.663
(0.005) (0.019) (0.024) (0.029) (0.013) (0.018) (0.020) (0.018) (0)
Strong regularity 0.649 0.983 0.986 0.984 0.624 0.981 0.981 0.984 0.999
(0.175) (0.006) (0.013) (0.008) (0.12) (0.015) (0.013) (0.011) (0)
Weak regularity 0.718 0.999 0.995 0.998 0.657 0.999 0.997 0.996 1.000
(0.165) (0.002) (0.007) (0.003) (0.125) (0.001) (0.003) (0.004) (0)
Panel 2: London dataset, sum-XGR
Log-likelihood $\mathbf{-506.5}$ $-542.9$ $-517.8$ $-560.6$ $\underline{-548.0}$ $-553.0$ $-565.9$ $-553.5$ $-570.2$
(4.002) (11.91) (6.636) (13.342) (3.359) (3.389) (5.310) (3.395) (0)
Accuracy 0.732 0.707 0.726 0.693 0.705 0.703 0.691 0.702 0.709
(0.005) (0.01) (0.006) (0.02) (0.005) (0.006) (0.007) (0.005) (0)
$F_{1}$ score 0.727 0.694 0.720 0.673 0.691 0.689 0.672 0.688 0.701
(0.005) (0.014) (0.007) (0.032) (0.006) (0.007) (0.008) (0.006) (0)
Strong regularity 0.898 0.999 0.991 0.999 0.946 0.968 0.998 0.970 1.000
(0.072) (0.002) (0.015) (0.003) (0.023) (0.013) (0.002) (0.009) (0)
Weak regularity 0.906 1.000 0.994 0.999 0.950 0.988 0.999 0.986 1.000
(0.072) (0.001) (0.012) (0.003) (0.021) (0.011) (0.001) (0.01) (0)
Table C.4: Model performance in the validation sets (1K-Random).
Panel 1: Chicago dataset, sum-XGR
@@
mean (std.) No GR PGR UGR LGR No GR PGR UGR LGR MNL
Log-likelihood $-151.9$ $\mathbf{-147.1}$ $-149.2$ $-147.6$ $-159.2$ $\underline{-155.3}$ $-155.5$ $-155.6$ $-163.0$
(1.033) (1.993) (1.876) (2.548) (1.725) (2.055) (1.920) (2.188) (0)
Accuracy 0.680 0.692 0.677 0.686 0.680 0.686 0.682 0.684 0.700
(0.006) (0.008) (0.008) (0.008) (0.011) (0.012) (0.012) (0.014) (0)
$F_{1}$ score 0.628 0.638 0.561 0.610 0.599 0.574 0.579 0.571 0.654
(0.006) (0.018) (0.016) (0.024) (0.019) (0.027) (0.026) (0.030) (0)
Strong regularity 0.671 0.984 0.990 0.988 0.692 0.982 0.980 0.985 0.995
(0.169) (0.009) (0.015) (0.015) (0.119) (0.018) (0.010) (0.012) (0)
Weak regularity 0.746 1.000 0.996 0.999 0.721 0.997 0.996 0.996 1.000
(0.154) (0.002) (0.007) (0.003) (0.119) (0.003) (0.005) (0.004) (0)
Panel 2: London dataset, sum-XGR
Log-likelihood $-143.3$ $\underline{-138.9}$ $-139.2$ $-140.5$ $-144.5$ $\underline{-141.7}$ $-143.4$ $-141.9$ $\mathbf{-138.7}$
(2.016) (1.676) (1.338) (3.175) (0.741) (0.958) (1.105) (1.085) (0)
Accuracy 0.688 0.681 0.693 0.680 0.697 0.705 0.698 0.705 0.700
(0.008) (0.015) (0.015) (0.019) (0.009) (0.009) (0.01) (0.009) (0)
$F_{1}$ score 0.683 0.663 0.685 0.653 0.676 0.684 0.676 0.684 0.689
(0.009) (0.019) (0.015) (0.031) (0.009) (0.01) (0.011) (0.01) (0)
Strong regularity 0.890 0.998 0.988 0.997 0.949 0.970 0.998 0.972 1.000
(0.073) (0.003) (0.021) (0.007) (0.027) (0.015) (0.003) (0.007) (0)
Weak regularity 0.896 0.999 0.992 0.997 0.951 0.984 0.998 0.982 1.000
(0.073) (0.002) (0.016) (0.006) (0.025) (0.012) (0.003) (0.012) (0)
Out-of-domain generalization
Table C.5: Model performance in the training sets (10K-Sorted).
Panel 1: Chicago dataset, sum-XGR
Metric: DNN TasteNet RUM
mean (std.) No GR PGR UGR LGR No GR PGR UGR LGR MNL
Log-likelihood $\mathbf{-4292.7}$ $-4325.0$ $-4398.3$ $-4350.1$ $-4847.9$ $-4600.0$ $-4595.5$ $\underline{-4590.9}$ $-4720.0$
(21.52) (17.68) (13.50) (20.77) (28.64) (9.302) (9.634) (11.26) (0)
Accuracy 0.745 0.742 0.738 0.741 0.715 0.726 0.727 0.727 0.726
(0.002) (0.002) (0.002) (0.002) (0.001) (0.001) (0.001) (0.001) (0)
$F_{1}$ score 0.717 0.714 0.707 0.713 0.661 0.691 0.692 0.692 0.693
(0.003) (0.003) (0.002) (0.003) (0.002) (0.002) (0.002) (0.001) (0)
Strong regularity 0.881 0.998 0.998 0.998 1.000 1.000 0.999 0.999 1.000
(0.045) (0.001) (0.001) (0.001) (0.000) (0.001) (0.001) (0.001) (0)
Weak regularity 0.898 1.000 1.000 1.000 1.000 1.000 1.000 0.999 1.000
(0.042) (0.000) (0.000) (0.000) (0.000) (0.001) (0.001) (0.001) (0)
Table C.5 (continued).
Panel 2: London dataset, sum-XGR
Metric: DNN TasteNet RUM
mean (std.) No GR PGR UGR LGR No GR PGR UGR LGR MNL
Log-likelihood $-4554.5$ $-4589.4$ $-4806.0$ $-4530.7$ $-4661.6$ $-4785.8$ $-4724.6$ $-4676.5$ $-4919.2$
(11.10) (25.82) (71.16) (6.477) (7.997) (41.15) (21.08) (8.192) (0)
Accuracy 0.725 0.722 0.710 0.725 0.724 0.712 0.717 0.721 0.712
(0.002) (0.002) (0.005) (0.002) (0.001) (0.005) (0.003) (0.001) (0)
$F_{1}$ score 0.722 0.719 0.706 0.722 0.720 0.706 0.713 0.717 0.708
(0.002) (0.002) (0.006) (0.002) (0.001) (0.006) (0.003) (0.001) (0)
Strong regularity 0.736 0.999 0.991 0.991 0.645 0.991 0.991 0.968 0.997
(0.06) (0.001) (0.019) (0.005) (0.07) (0.008) (0.006) (0.023) (0)
Weak regularity 0.751 0.999 0.993 0.994 0.676 1.000 0.999 0.989 1.000
(0.054) (0.0) (0.016) (0.002) (0.067) (0.000) (0.001) (0.005) (0)
Table C.6: Model performance in the validation sets (10K-Sorted).
Panel 1: Chicago dataset, sum-XGR
Metric: DNN TasteNet RUM
mean (std.) No GR PGR UGR LGR No GR PGR UGR LGR MNL
Log-likelihood $-656.6$ $-647.9$ $-648.5$ $\mathbf{-646.9}$ $-683.0$ $-660.1$ $\underline{-660.0}$ $-660.2$ $-673.2$
(4.636) (3.430) (2.176) (2.397) (3.344) (1.171) (1.316) (1.320) (0)
Accuracy 0.737 0.739 0.737 0.739 0.729 0.745 0.746 0.745 0.745
(0.003) (0.002) (0.004) (0.003) (0.004) (0.003) (0.002) (0.002) (0)
$F_{1}$ score 0.703 0.707 0.704 0.707 0.675 0.709 0.710 0.709 0.711
(0.004) (0.002) (0.005) (0.005) (0.005) (0.004) (0.003) (0.003) (0)
Strong regularity 0.883 0.998 0.998 0.998 1.000 1.000 1.000 1.000 1.000
(0.045) (0.001) (0.001) (0.001) (0.000) (0.001) (0.001) (0.001) (0)
Weak regularity 0.898 0.999 1.000 1.000 1.000 1.000 1.000 1.000 1.000
(0.042) (0.001) (0.000) (0.000) (0.000) (0.001) (0.001) (0.001) (0)
Panel 2: London dataset, sum-XGR
Log-likelihood $-698.5$ $-699.7$ $-713.6$ $\mathbf{-696.3}$ $-720.3$ $\underline{-726.6}$ $-722.7$ $-720.4$ $-730.2$
(2.201) (2.208) (4.791) (2.877) (2.160) (4.988) (3.051) (2.274) (0)
Accuracy 0.691 0.693 0.691 0.694 0.701 0.686 0.694 0.699 0.687
(0.004) (0.005) (0.004) (0.004) (0.003) (0.008) (0.005) (0.003) (0)
F1 score 0.688 0.690 0.688 0.692 0.697 0.680 0.689 0.695 0.683
(0.004) (0.005) (0.004) (0.004) (0.003) (0.009) (0.005) (0.003) (0)
Strong regularity 0.724 0.998 0.991 0.990 0.637 0.992 0.992 0.967 0.996
(0.062) (0.001) (0.020) (0.004) (0.070) (0.009) (0.006) (0.026) (0)
Weak regularity 0.739 0.999 0.992 0.992 0.668 1.000 0.999 0.988 1.000
(0.054) (0.001) (0.018) (0.003) (0.066) (0.001) (0.001) (0.007) (0)
Appendix D Individual demands as functions of regularization strength

### v2

@@ Appendix C Training and validation performance
Mean (SD) No GR PGR UGR LGR No GR PGR UGR LGR ReLU Exp MNL
Panel 1: CMAP data, sum-XGR
Log-likelihood $\mathbf{-494.8}$ $\underline{-511.0}$ $-546.4$ $-526.2$ $-555.6$ $-580.2$ $-570.9$ $-583.1$ $-563.0$ $-558.1$ $-548.1$
(6.074) (8.266) (9.391) (8.871) (4.071) (6.789) (5.316) (9.196) (4.542) (6.232) (0)
Accuracy 0.740 0.732 0.691 0.718 0.711 0.691 0.696 0.690 0.708 0.718 0.715
(0.005) (0.011) (0.008) (0.01) (0.006) (0.007) (0.008) (0.007) (0.007) (0.006) (0)
$F_{1}$ score 0.690 0.682 0.578 0.648 0.634 0.575 0.588 0.573 0.623 0.652 0.663
(0.005) (0.019) (0.024) (0.029) (0.013) (0.018) (0.020) (0.018) (0.012) (0.008) (0)
Strong regularity 0.649 0.983 0.986 0.984 0.624 0.981 0.981 0.984 0.450 1.000 0.999
(0.175) (0.006) (0.013) (0.008) (0.120) (0.015) (0.013) (0.011) (0.185) (0.000) (0)
Weak regularity 0.718 0.999 0.995 0.998 0.657 0.999 0.997 0.996 1.000 1.000 1.000
(0.165) (0.002) (0.007) (0.003) (0.125) (0.001) (0.003) (0.004) (0.000) (0.000) (0)
Panel 2: LTDS data, sum-XGR
Log-likelihood $\mathbf{-506.5}$ $-542.9$ $\underline{-517.8}$ $-560.6$ $-548.0$ $-553.0$ $-565.9$ $-553.5$ $-562.1$ $-541.0$ $-570.2$
(4.002) (11.91) (6.636) (13.342) (3.359) (3.389) (5.310) (3.395) (10.07) (3.794) (0)
Accuracy 0.732 0.707 0.726 0.693 0.705 0.703 0.691 0.702 0.703 0.720 0.709
(0.005) (0.010) (0.006) (0.02) (0.005) (0.006) (0.007) (0.005) (0.007) (0.007) (0)
$F_{1}$ score 0.727 0.694 0.720 0.673 0.691 0.689 0.672 0.688 0.690 0.714 0.701
(0.005) (0.014) (0.007) (0.032) (0.006) (0.007) (0.008) (0.006) (0.007) (0.007) (0)
Strong regularity 0.898 0.999 0.991 0.999 0.946 0.968 0.998 0.970 0.931 0.995 1.000
(0.072) (0.002) (0.015) (0.003) (0.023) (0.013) (0.002) (0.009) (0.047) (0.003) (0)
Weak regularity 0.906 1.000 0.994 0.999 0.950 0.988 0.999 0.986 1.000 1.000 1.000
(0.072) (0.001) (0.012) (0.003) (0.021) (0.011) (0.001) (0.01) (0.000) (0.000) (0)
Table 14: Model performance in the validation sets of 1K-Random.
@@
Mean (SD) No GR PGR UGR LGR No GR PGR UGR LGR ReLU Exp MNL
Panel 1: CMAP data, sum-XGR
Log-likelihood $-151.9$ $\mathbf{-147.1}$ $-149.2$ $\underline{-147.6}$ $-159.2$ $-155.3$ $-155.5$ $-155.6$ $-157.8$ $-160.4$ $-163.0$
(1.033) (1.993) (1.876) (2.548) (1.725) (2.055) (1.920) (2.188) (2.613) (2.203) (0)
Accuracy 0.680 0.692 0.677 0.686 0.680 0.686 0.682 0.684 0.692 0.697 0.700
(0.006) (0.008) (0.008) (0.008) (0.011) (0.012) (0.012) (0.014) (0.010) (0.012) (0)
$F_{1}$ score 0.628 0.638 0.561 0.610 0.599 0.574 0.579 0.571 0.606 0.637 0.654
(0.006) (0.018) (0.016) (0.024) (0.019) (0.027) (0.026) (0.030) (0.017) (0.017) (0)
Strong regularity 0.671 0.984 0.990 0.988 0.692 0.982 0.980 0.985 0.468 0.994 0.995
(0.169) (0.009) (0.015) (0.015) (0.119) (0.018) (0.010) (0.012) (0.208) (0.003) (0)
Weak regularity 0.746 1.000 0.996 0.999 0.721 0.997 0.996 0.996 1.000 1.000 1.000
(0.154) (0.002) (0.007) (0.003) (0.119) (0.003) (0.005) (0.004) (0.000) (0.000) (0)
Panel 2: LTDS data, sum-XGR
Log-likelihood $-143.3$ $\underline{-138.9}$ $-139.2$ $-140.5$ $-144.5$ $-141.7$ $-143.4$ $-141.9$ $-145.3$ $-139.1$ $\mathbf{-138.7}$
(2.016) (1.676) (1.338) (3.175) (0.741) (0.958) (1.105) (1.085) (2.642) (1.596) (0)
Accuracy 0.688 0.681 0.693 0.680 0.697 0.705 0.698 0.705 0.694 0.727 0.700
(0.008) (0.015) (0.015) (0.019) (0.009) (0.009) (0.010) (0.009) (0.013) (0.014) (0)
$F_{1}$ score 0.683 0.663 0.685 0.653 0.676 0.684 0.676 0.684 0.674 0.719 0.689
(0.009) (0.019) (0.015) (0.031) (0.009) (0.010) (0.011) (0.01) (0.013) (0.014) (0)
Strong regularity 0.890 0.998 0.988 0.997 0.949 0.970 0.998 0.972 0.937 1.000 1.000
(0.073) (0.003) (0.021) (0.007) (0.027) (0.015) (0.003) (0.007) (0.044) (0.000) (0)
Weak regularity 0.896 0.999 0.992 0.997 0.951 0.984 0.998 0.982 1.000 1.000 1.000
(0.073) (0.002) (0.016) (0.006) (0.025) (0.012) (0.003) (0.012) (0.000) (0.000) (0)
Out-of-domain generalization
See Tables 15 and 16.
Table 15: Model performance in the training sets of 10K-Sorted.
@@
Mean (SD) No GR PGR UGR LGR No GR PGR UGR LGR ReLU Exp MNL
Panel 1: CMAP data, sum-XGR
Log-likelihood $\mathbf{-4292.7}$ $\underline{-4325.0}$ $-4398.3$ $-4350.1$ $-4847.9$ $-4600.0$ $-4595.5$ $-4590.9$ $-4664.2$ $-4572.7$ $-4720.0$
(21.52) (17.68) (13.50) (20.77) (28.64) (9.302) (9.634) (11.26) (60.61) (8.131) (0)
Accuracy 0.745 0.742 0.738 0.741 0.715 0.726 0.727 0.727 0.723 0.730 0.726
(0.002) (0.002) (0.002) (0.002) (0.001) (0.001) (0.001) (0.001) (0.003) (0.001) (0)
$F_{1}$ score 0.717 0.714 0.707 0.713 0.661 0.691 0.692 0.692 0.684 0.698 0.693
(0.003) (0.003) (0.002) (0.003) (0.002) (0.002) (0.002) (0.001) (0.006) (0.001) (0)
Strong regularity 0.881 0.998 0.998 0.998 1.000 1.000 0.999 0.999 0.952 1.000 1.000
(0.045) (0.001) (0.001) (0.001) (0.000) (0.001) (0.001) (0.001) (0.141) (0.000) (0)
Weak regularity 0.898 1.000 1.000 1.000 1.000 1.000 1.000 0.999 1.000 1.000 1.000
(0.042) (0.000) (0.000) (0.000) (0.000) (0.001) (0.001) (0.001) (0.000) (0.000) (0)
Panel 2: LTDS data, sum-XGR
Log-likelihood $\underline{-4554.5}$ $-4589.4$ $-4806.0$ $\mathbf{-4530.7}$ $-4661.6$ $-4785.8$ $-4724.6$ $-4676.5$ $-4752.8$ $-4663.7$ $-4919.2$
(11.10) (25.82) (71.16) (6.477) (7.997) (41.15) (21.08) (8.192) (75.76) (8.416) (0)
Accuracy 0.725 0.722 0.710 0.725 0.724 0.712 0.717 0.721 0.716 0.724 0.712
(0.002) (0.002) (0.005) (0.002) (0.001) (0.005) (0.003) (0.001) (0.004) (0.001) (0)
$F_{1}$ score 0.722 0.719 0.706 0.722 0.720 0.706 0.713 0.717 0.712 0.720 0.708
(0.002) (0.002) (0.006) (0.002) (0.001) (0.006) (0.003) (0.001) (0.005) (0.001) (0)
Strong regularity 0.736 0.999 0.991 0.991 0.645 0.991 0.991 0.968 0.430 0.998 0.997
(0.06) (0.001) (0.019) (0.005) (0.07) (0.008) (0.006) (0.023) (0.172) (0.000) (0)
Weak regularity 0.751 0.999 0.993 0.994 0.676 1.000 0.999 0.989 1.000 1.000 1.000
(0.054) (0.0) (0.016) (0.002) (0.067) (0.000) (0.001) (0.005) (0.000) (0.000) (0)
Table 16: Model performance in the validation sets of 10K-Sorted.
@@
Mean (SD) No GR PGR UGR LGR No GR PGR UGR LGR ReLU Exp MNL
Panel 1: CMAP data, sum-XGR
Log-likelihood $-656.6$ $\underline{-647.9}$ $-648.5$ $\mathbf{-646.9}$ $-683.0$ $-660.1$ $-660.0$ $-660.2$ $-665.1$ $-671.2$ $-673.2$
(4.636) (3.430) (2.176) (2.397) (3.344) (1.171) (1.316) (1.320) (5.144) (1.098) (0)
Accuracy 0.737 0.739 0.737 0.739 0.729 0.745 0.746 0.745 0.744 0.741 0.745
(0.003) (0.002) (0.004) (0.003) (0.004) (0.003) (0.002) (0.002) (0.003) (0.004) (0)
$F_{1}$ score 0.703 0.707 0.704 0.707 0.675 0.709 0.710 0.709 0.704 0.707 0.711
(0.004) (0.002) (0.005) (0.005) (0.005) (0.004) (0.003) (0.003) (0.004) (0.005) (0)
Strong regularity 0.883 0.998 0.998 0.998 1.000 1.000 1.000 1.000 0.952 0.999 1.000
(0.045) (0.001) (0.001) (0.001) (0.000) (0.001) (0.001) (0.001) (0.142) (0.000) (0)
Weak regularity 0.898 0.999 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
(0.042) (0.001) (0.000) (0.000) (0.000) (0.001) (0.001) (0.001) (0.000) (0.000) (0)
Panel 2: LTDS data, sum-XGR
Log-likelihood $\underline{-698.5}$ $-699.7$ $-713.6$ $\mathbf{-696.3}$ $-720.3$ $-726.6$ $-722.7$ $-720.4$ $-721.9$ $-719.4$ $-730.2$
(2.201) (2.208) (4.791) (2.877) (2.160) (4.988) (3.051) (2.274) (2.608) (1.861) (0)
Accuracy 0.691 0.693 0.691 0.694 0.701 0.686 0.694 0.699 0.698 0.697 0.687
(0.004) (0.005) (0.004) (0.004) (0.003) (0.008) (0.005) (0.003) (0.004) (0.002) (0)
F1 score 0.688 0.690 0.688 0.692 0.697 0.680 0.689 0.695 0.693 0.693 0.683
(0.004) (0.005) (0.004) (0.004) (0.003) (0.009) (0.005) (0.003) (0.004) (0.002) (0)
Strong regularity 0.724 0.998 0.991 0.990 0.637 0.992 0.992 0.967 0.420 0.997 0.996
(0.062) (0.001) (0.020) (0.004) (0.070) (0.009) (0.006) (0.026) (0.168) (0.001) (0)
Weak regularity 0.739 0.999 0.992 0.992 0.668 1.000 0.999 0.988 1.000 1.000 1.000
(0.054) (0.001) (0.018) (0.003) (0.066) (0.001) (0.001) (0.007) (0.000) (0.000) (0)
Appendix D Effects of regularization strength
Fig. 10 visualizes the effects of regularization strength $\lambda$ on individual demand functions.